# TiDES: The 4MOST Time Domain Extragalactic Survey

### v1

@@ 3.5 LSST-TiDES selection function
We next describe the process of selecting transients from the LSST simulations for 4MOST spectroscopic follow-up. The aim is to reflect the ‘real-life’ operation of TiDES, where live SNe are identified through difference imaging and distributed in alert packets in the LSST data stream, we will identify suitable targets through custom filtering in community full-stream alert brokers (e.g., Lasair, ALeRCE, ANTARES, AMPEL, FINK, Pitt-Google, Babmul; Smith et al., 2019; Förster et al., 2021; Matheson et al., 2021; Nordin et al., 2019; Möller et al., 2021; Wood-Vasey, 2024). This means that new targets must be frequently added to the TiDES target list, and targets that have faded must be removed.
The depth of LSST is fainter than the spectroscopic limiting magnitude of 4MOST, allowing us to track the rising brightness of a transient before submitting it to the 4MOST observing queue – resulting in a high purity of candidates. Significant advances have been made in rapid real-time classification of transients and spectroscopic follow-up recommendation algorithms (e.g., Muthukrishna et al., 2019a; Kennamer et al., 2020; Qu & Sako, 2022; Gagliano et al., 2023; Möller et al., 2024; Sheng et al., 2024). However, TiDES has access to sufficient fibres to observe all live transients in any field down to its limiting magnitude, and so is able to adopt a very straight forward selection of potential targets. This also has the advantage of a simpler calculation of the efficiency and selection function for any subsequent TiDES population analysis, including for SN Ia cosmology.
@@
Once all criteria are met, the target is added to a mock TiDES observing queue with a nominal lifetime of four days, after which the target is retired. Each object, however, is monitored so that each subsequent LSST observation of the target that is brighter than 22.5 mag refreshes the target in the queue by extending its expiry date by an additional four days. This process ensures a high probability that any target selected by the 4MOST scheduler will be bright enough to have a successful spectroscopic observation (see our spectral success criteria in Section 4.2).
@@ 4 Expected results from 4MOST/TiDES
The results from our simulations provide us with a catalogue of SNe, their hosts, and AGN that were ‘observed’ as part of the survey realisation. This includes metadata for the observational conditions, including the seeing, exposure time, and airmass. Combining all of this, we are able to generate mock spectra for each of our observed events, evaluate the quality of the mock spectra, and then quantify the performance of the survey simulation. We note that the conclusions we draw here are demonstrated at the population-level, and we do not consider any individual object assessment or seek to derive astrophysical properties from the spectra.
@@ 4.1 Mock spectra
To generate mock spectra, template spectra for each observed event are used as inputs to the 4MOST exposure-time calculator444https://escience.aip.de/readthedocs/OpSys/etc/master/index.html, with parameters matching the observing conditions from the simulations. For the TiDES-Live survey, we generate phase-specific spectra based on the date of the 4MOST observation using the same templates described in 3.1. Where the same SN is observed more than once, we do not stack the spectra to improve the signal-to-noise due to evolution in the target spectrum. Conversely, for the TiDES-Hosts our targets can be repeatedly observed on multiple occasions and are not expected to evolve, and we take the total integrated exposure time over the survey when generating our final spectra. The ETC returns the instrument-level signal, noise, sky background, and the efficiencies for each of the three arms on the LRS.
@@ 4.2 Spectral success criteria
For TiDES-Live, our actual success will depend on the signal-to-noise ratio (SNR) in the 4MOST spectrum and the transient type, due to the prominence of certain spectral features that define the different classes. SN spectra are typically dominated by broad features in the photospheric phases that are many tens-of-Angstroms wide, and thus our SSC are defined using the mean SNR in 15Å bins – we denote this with the SNR parameter. Our criteria are based on earlier studies of high-redshift SNe Ia (Balland et al., 2009), where a robust classification was achieved with a mean SNR = 5 over 4500–8000Å in the observer frame. This is a conservative criterion as Balland et al. (2009) also demonstrate probable classifications of transients with a mean SNR = 3.
We have checked our likely classification limits for SNe Ia using the ETC based on simulated mock spectra with classifications based upon the machine-learning SN classification tool dash (Muthukrishna et al., 2019b). dash can classify, with 95 per cent confidence, a SN Ia spectrum with SNR = 5 without using the host redshift as a prior. These simulations were performed with SN spectra free of host galaxy contamination (e.g., the SN Ia spectra in Fig. 3). An investigation into the effect of host galaxy contamination on SN classification in 4MOST will be presented in Milligan et al.(in prep).
@@
For TiDES-RM, spectral success is defined by achieving a SNR =10 for an AGN spectrum. Just as for the TiDES-Live program, the spectra vary with time and, therefore, cannot be stacked if observations are taken on significantly different epochs. Each AGN in the sample is observed with a cadence of d under a special 4MOST ‘mini-survey’ that will be conducted on the LSST DDFs. To meet the SSC, the TiDES-RM targets must reach SNR =10 in a single visit every 14 d.
@@ 4.3 Live supernova populations
We begin with the most general overview metric for any wide-field sky-survey – the redshift distribution. Fig. 4 shows the redshift distribution split into the different transient sub-samples from our simulations. Our results show that we can expect to observe around 30 000 SNe with (SNR ) in total. Due to their intrinsic brightness, SNe Ia are the most common transient that TiDES will observe, totalling around 18 000 objects with a peak in the redshift distribution at z. This results in the largest homogeneously collected SN Ia sample from any spectroscopic survey. As expected, the SN1991bg-like ( and SN Iax under-luminous populations of SNe Ia are found at lower redshifts and are intrinsically rarer than the normal counterparts.
SNe II are intrinsically the most common of all SN sub-types (e.g., Li et al., 2011; Graur et al., 2017b) and this is reflected in the first two redshift bins of the histogram in Fig. 4. However, their lower luminosity results in a lower peak redshift for the sample at z. Stripped-envelope SNe (SESNe) also belong to the massive-star progenitor, core-collapse SN family, but occur at per cent of the intrinsic rate of SNe II (Frohmaier et al., 2021). In our sample, SESNe follow a similar observed trend with a peak redshift of z.
SLSNe-I form a class of the brightest observed SNe with a luminosity function ranging from (Angus et al., 2019). As a result, they are observable to the highest redshifts of any objects in our simulated sample, and their intrinsically long (and time dilated) light curves increases the opportunities for TiDES to obtain a spectrum. We have truncated the SLSN redshift distribution in Fig. 4 for visualisation purposes, but note that we expect to obtain SLSNe-I up to z.
Given the intrinsic luminosity, rates, and light curve evolution timescales of both Calcium-strong transients (CaST) and TDEs, it is unsurprising that these objects form the smallest populations of our samples. Furthermore, as these are some of the least well understood transient classes, any recent updates to their observed properties will not be reflected in our transient models from PLAsTiCC. For example a slightly brighter TDE luminosity function than the one we adopt is presented in Yao et al. (2023). Regardless, the TiDES sample of these transients will be a significant increase over current sample sizes.
We analyse the efficiency of our survey as a function of the target object’s magnitude in Fig. 5. We define the efficiency as the fraction of objects meeting the SSC once a fibre has been placed on them. To estimate these efficiencies, we fit the following sigmoid of the functional form (Sharon et al., 2007; Rodney et al., 2014):
where is the efficiency, is the maximum efficiency, is the magnitude of the object in the LSST -band, is the magnitude at which 50 per cent of the sample meet our SSC, and captures the exponential roll-off (smaller numbers describe a steeper drop).
@@ 4.5 TiDES host galaxies
TiDES-Hosts is focused on spectroscopy of SN host galaxies. At the most fundamental level we aim to obtain secure redshifts for these hosts but, as exposure depth is built up for some targets, the increasing signal-to-noise will enable us to measure additional properties (e.g., star formation rate, gas phase metallicity (Dumayne et al., 2023)).
@@ 5 Supernova Cosmology
Following the forecasts presented in the Dark Energy Science Collaboration (DESC) Science Requirement Document (The LSST Dark Energy Science Collaboration et al., 2018), we also include an external low- sample in our cosmology analysis of 2 400 SNe Ia at . The low- SNe in our analysis follows the simulations of the Foundation SN sample (Foley et al., 2018) presented by Jones et al. (2019). The expectation, however, is that this final sample size will only be achieved towards the end of 10-years of LSST operations, even though TiDES is expected to conclude after the first five years of LSST.
@@ 5.1 Defining the SN sample
We fit all simulated SNe Ia with the SALT2 light curve fitter (Guy et al., 2007) based on the simulated observed photometry. To ensure meaningful light curve fits, we follow standard requirements (e.g. Brout et al., 2019) of: i) observations in at least two different filters with SNR5, ii) at least one data point before the time of peak brightness, , and iii) at least one data point ten days after . The first of these are already satisfied by the the spectroscopic triggering criteria in Section 3.5, and the final two are satisfied in the majority of cases due to the high cadence of WFD in LSST.
Once the model has been fit to light curves, a final subset of the SNe Ia is selected by applying the following SALT2-based cuts (e.g., Betoule et al., 2014; Brout et al., 2019): , , , days, and fit probability 0.001. Those that pass this step are considered ‘cosmologically useful’ supernovae. These selection criteria are generally adopted in SN Ia cosmology analyses to reduce contamination from highly-reddened or other peculiar thermonuclear SNe while ensuring a minimal loss of ‘cosmologically useful’ SN Ia events. We find 12 600 successful SALT2 fits in the spectroscopic sample and 131 000 successful fits in the photometric SN sample. The redshift distribution for both samples and the low- sample are shown in the top panel of Fig. 11.
@@ 5.2 SN distance estimation and bias corrections
After light-curve fitting, we measure SN Ia distances and build the redshift–distance relation (‘Hubble diagram’) from which cosmological parameters can be constrained. We estimate the SN distance modulus , which is proportional to the logarithm of the SN luminosity distance, by applying the Tripp formula (Tripp, 1998)
where symbols are as defined in Section 3.1.1. The nuisance parameters , and are determined following the approach presented in Marriner et al. (2011). We present our final Hubble diagram in Fig. 11.
The term in Equation 3 is a bias correction term that is applied to each SN to correct for survey selection effects. In this work, bias corrections are measured using the BEAMS with Bias Corrections (BBC; Kessler et al., 2019b) code. BBC uses large simulations of the survey to model and correct for selection effects while simultaneously fitting for the nuisance parameters and intrinsic SN scatter. For our analysis, corrections for selection effects are modelled as a function of redshift only (so-called ‘1D’ corrections) and they are estimated using the same LSST+TiDES simulated samples used to measure the cosmological constraints.
The output of BBC is a redshift-binned Hubble diagram corrected for selection effects, and the associated diagonal covariance matrix, , that includes statistical uncertainties only for each redshift bin. In the next Section, we describe how this systematic covariance matrix is built.
@@ 5.3 Cosmological Parameters
The cosmological contours from our simulations are presented in Fig. 12 and Fig. 13. We present constraints from the LSST+TiDES SN events (both spectroscopically-confirmed SNe Ia and SNe Ia with host spec- only) and the external low- sample. We compare our results with constraints from a simulation of the Dark Energy Survey 5 year SN sample (DES-SN5YR, The DES Collaboration, 2024). The DES-SN5YR analysis constitutes the state-of-the-art of SN cosmological measurements and provides some of the tightest constraints on the Dark Energy Equation of State. The DES-SN5YR simulation used in this work is generating assuming the same cosmology used in our LSST simulations (i.e., FlatCDM cosmology with ) and reflects the statistical power and data quality of the real DES-SN5YR data. While the published DES-SN5YR include statistical and systematic uncertainties, in Fig. 12 and Fig. 13, we only consider statistical uncertainties. For a FlatCDM model, LSST+TiDES SN samples combined with external low- samples provide constrains on the Dark Energy Equation of State parameter, , of 0.012, 10 times smaller compared to DES-SN5YR. For a FlatCDM model, the Figure of Merit obtained from LSST+TiDES SN samples combined with external low- samples is 85, 15 times larger than the FoM of DES-SN5YR. When fitting for a FlatCDM model, we also combine SN data with a prior based on the recent Baryonic Acoustic Oscillation (BAO) measurements published by the Dark Energy Spectroscopic Instrument (DESI) Collaboration (DESI Collaboration et al., 2024) using their first year of data. The DESI-BAO-Y1 prior is applied assuming the same true cosmology used to generate the LSST and DES-SN5YR simulations. With the DESI-BAO-Y1 prior, the LSST+TiDES FoM increases to 99 (7 times larger compared to DES-SN5YR combined with the same DESI-BAO-Y1 prior).
@@ 6 Summary
We performed a cosmology analysis of the LSST+TiDES sample, including a low- sample predicted by Jones et al. (2019). We are able to constrain the dark energy equations-of-state parameter, , to a sub-2 per cent statistical-only precision, this is 10 times smaller compared to the current gold-standard (DES Collaboration et al., 2024).
@@
Software
Acknowledgements
The authors would like to thank Ricardo Demarco and Swayamtrupta Panda for their comments on this manuscript.
This paper has undergone internal review in the LSST Dark Energy Science Collaboration. The internal reviewers were Chris Lidman and Alex Gagliano.
CF and MSu acknowledge support from STFC funding for UK participation in LSST, through grants ST/V002031/1 and ST/X00130X/1.
MN is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 948381) and by UK Space Agency Grant No. ST/Y000692/1.
SFH acknowledges funding from STFC Grant ST/Y001656/1 and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding Guarantee (EP/Z533920/1).
SJS acknowledges funding from STFC Grants ST/Y001605/1, ST/X001253/1, ST/X006506/1, ST/T000198/1 and the a Royal Society Research Professorship grant.
AM is supported by the ARC Discovery Early Career Researcher Award (DECRA) project number DE230100055.
Y.-L.K. has received funding from the Science and Technology Facilities Council [grant number ST/V000713/1].
LK acknowledges support for an Early Career Fellowship from the Leverhulme Trust through grant ECF-2024-054 and the Isaac Newton Trust through grant 24.08(w).
ET acknowledges funding from the HTM (grant TK202), ETAg (grant PRG1006) and the EU Horizon Europe (EXCOSM, grant No. 101159513).
CF led the analysis, performed the simulations, and wrote the paper. MV and MSm performed the cosmology analysis. MSu discussed the results, implications, and commented on the manuscript at all stages of the project. SFH is PI of the TiDES-RM program and wrote the relevant sections of this paper.
The DESC acknowledges ongoing support from the Institut National de Physique Nucléaire et de Physique des Particules in France; the Science & Technology Facilities Council in the United Kingdom; and the Department of Energy, the National Science Foundation, and the LSST Corporation in the United States. DESC uses resources of the IN2P3 Computing Center (CC-IN2P3–Lyon/Villeurbanne - France) funded by the Centre National de la Recherche Scientifique; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231; STFC DiRAC HPC Facilities, funded by UK BEIS National E-infrastructure capital grants; and the UK particle physics grid, supported by the GridPP Collaboration. This work was performed in part under DOE Contract DE-AC02-76SF00515.

### v2

@@ 3.5 LSST-TiDES selection function
We next describe the process of selecting transients from the LSST simulations for 4MOST spectroscopic follow-up. The aim is to reflect the ‘real-life’ operation of TiDES, where live SNe are identified through difference imaging and distributed in alert packets in the LSST data stream, we will identify suitable targets through custom filtering in community full-stream alert brokers e.g. Lasair (K. W. Smith et al., 2019), ALeRCE (F. Förster et al., 2021), ANTARES (T. Matheson et al., 2021), AMPEL (J. Nordin et al., 2019), FINK (A. Möller et al., 2021), Pitt-Google (W. M. Wood-Vasey, 2024), and Babamul. This means that new targets must be frequently added to the TiDES target list, and targets that have faded must be removed.
The depth of LSST is fainter than the spectroscopic limiting magnitude of 4MOST, allowing us to track the rising brightness of a transient before submitting it to the 4MOST observing queue – resulting in a high purity of candidates. Significant advances have been made in rapid real-time classification of transients and spectroscopic follow-up recommendation algorithms (e.g., D. Muthukrishna et al., 2019a; N. Kennamer et al., 2020; H. Qu & M. Sako, 2022; A. Gagliano et al., 2023; A. Möller et al., 2024; X. Sheng et al., 2024). However, TiDES has access to sufficient fibres to observe all live transients444In well-sampled LSST fields, live-transients in a 4MOST FoV pass our selection criteria in any field down to its limiting magnitude, and so is able to adopt a very straight forward selection of potential targets. This also has the advantage of a simpler calculation of the efficiency and selection function for any subsequent TiDES population analysis, including for SN Ia cosmology.
@@
Once all criteria are met, the target is added to a mock TiDES observing queue with a nominal lifetime of four days, after which the target is retired. Each object, however, is monitored so that each subsequent LSST observation of the target that is brighter than 22.5 mag refreshes the target in the queue by extending its expiry date by an additional four days. This process ensures a high probability that any target selected by the 4MOST scheduler will be bright enough to have a successful spectroscopic observation (see our Spectral Success Criteria (SSC) in Section 4.2).
@@ 4 Expected results from 4MOST/TiDES
The results from our simulations provide us with a catalogue of SNe, their hosts, and AGN that were ‘observed’ as part of the survey realisation. This includes metadata for the observational conditions, including the seeing, exposure time, and airmass. Combining all of this, we are able to generate mock spectra for each of our observed events, evaluate the quality of the mock spectra, and then quantify the performance of the survey simulation. We note that the conclusions we draw here are demonstrated at the population-level, and we do not consider any individual object assessment or seek to derive astrophysical properties from the spectra. Based on the density of targets and signal-to-noise requirements for each of our three science programmes, our fibre-hours are proportionally spent between TiDES-SN, TiDES-Hosts, and TiDES-RM respectively.
@@ 4.1 Mock spectra
To generate mock spectra, template spectra for each observed event are used as inputs to the 4MOST exposure-time calculator (ETC)555https://escience.aip.de/readthedocs/OpSys/etc/master/index.html, with parameters matching the observing conditions from the simulations. For the TiDES-Live survey, we generate phase-specific spectra based on the date of the 4MOST observation using the same templates described in 3.1. Where the same SN is observed more than once, we do not stack the spectra to improve the signal-to-noise due to evolution in the target spectrum. Conversely, for the TiDES-Hosts our targets can be repeatedly observed on multiple occasions and are not expected to evolve, and we take the total integrated exposure time over the survey when generating our final spectra. The ETC returns the instrument-level signal, noise, sky background, and the efficiencies for each of the three arms on the LRS.
@@ 4.2 Spectral success criteria
For TiDES-Live, our actual success will depend on the signal-to-noise ratio (SNR) in the 4MOST spectrum and the transient type, due to the prominence of certain spectral features that define the different classes. SN spectra are typically dominated by broad features in the photospheric phases that are many tens-of-Angstroms wide, and thus our SSC are defined using the mean SNR in 15Å bins – we denote this with the SNR parameter. Our criteria are based on earlier studies of high-redshift SNe Ia (C. Balland et al., 2009), where a robust classification was achieved with a mean SNR = 5 over 4500–8000Å in the observer frame. This is a conservative criterion as C. Balland et al. (2009) also demonstrate probable classifications of transients with a mean SNR = 3.
We have checked our likely classification limits for SNe Ia using the ETC based on simulated mock spectra with classifications based upon the machine-learning SN classification tool dash (D. Muthukrishna et al., 2019b). dash can classify, with 95 per cent confidence, a SN Ia spectrum with SNR = 5 without using the host redshift as a prior. These simulations were performed with SN spectra free of host galaxy contamination (e.g., the SN Ia spectra in Fig. 3). An investigation into the effect of host galaxy contamination on SN classification in 4MOST will be presented in Milligan et al.(in prep).
@@
For TiDES-RM, spectral success is defined by achieving a SNR =10 for an AGN spectrum. Just as for the TiDES-Live program, the spectra vary with time and, therefore, cannot be stacked if observations are taken on significantly different epochs. Within 4MOST, the other participating surveys have placed a high demand of time on the LSST DDFs. This necessitates many repeat visits that TiDES will exploit to routinely observe our AGN for a reverberation mapping experiment. To meet the SSC, the TiDES-RM targets must reach SNR =10 in a single visit every d.
@@ 4.3 Live supernova populations
We begin with the most general overview metric for any wide-field sky-survey – the redshift distribution. Fig. 4 shows the redshift distribution split into the different transient sub-samples from our simulations. Our results show that we can expect to observe around 30 000 SNe with (SNR ) in total. Due to their intrinsic brightness, SNe Ia are the most common transient that TiDES will observe, totalling around 18 000 objects with a peak in the redshift distribution at z. This results in the largest homogeneously collected SN Ia sample from any spectroscopic survey. As expected, the SN1991bg-like and SN Iax under-luminous populations of SNe Ia are found at lower redshifts and are intrinsically rarer than the normal counterparts.
SNe II are intrinsically the most common of all SN sub-types (e.g., W. Li et al., 2011; O. Graur et al., 2017b) and this is reflected in the first two redshift bins of the histogram in Fig. 4. However, their lower luminosity results in a lower peak redshift for the sample at z. Stripped-envelope SNe (SESNe) also belong to the massive-star progenitor, core-collapse SN family, but occur at per cent of the intrinsic rate of SNe II (C. Frohmaier et al., 2021). In our sample, SESNe follow a similar observed trend with a peak redshift of z.
SLSNe-I form a class of the brightest observed SNe with a luminosity function ranging from (C. R. Angus et al., 2019). As a result, they are observable to the highest redshifts of any objects in our simulated sample, and their intrinsically long (and time dilated) light curves increases the opportunities for TiDES to obtain a spectrum. We have truncated the SLSN redshift distribution in Fig. 4 for visualisation purposes, but note that we expect to obtain SLSNe-I up to z.
Given the intrinsic luminosity, rates, and light curve evolution timescales of both Calcium-strong transients (CaST) and TDEs, it is unsurprising that these objects form the smallest populations of our samples. Furthermore, as these are some of the least well understood transient classes, any recent updates to their observed properties will not be reflected in our transient models from PLAsTiCC. For example, a slightly brighter TDE luminosity function than the one we adopt is presented in Y. Yao et al. (2023). Regardless, the TiDES sample of these transients will be a significant increase over current sample sizes.
We analyse the efficiency of our survey as a function of the target object’s magnitude in Fig. 5. We define the efficiency as the fraction of objects meeting the SSC once a fibre has been placed on them. To estimate these efficiencies, we fit the following sigmoid of the functional form (K. Sharon et al., 2007; S. A. Rodney et al., 2014):
where is the efficiency, is the maximum efficiency, is the magnitude of the object in the LSST -band at the time of observation, is the magnitude at which 50 per cent of the sample meet our SSC, and captures the exponential roll-off (smaller numbers describe a steeper drop).
@@ 4.5 TiDES host galaxies
TiDES-Hosts is focused on spectroscopy of SN host galaxies. At the most fundamental level we aim to obtain secure redshifts for these hosts but, as exposure depth is built up for some targets, the increasing signal-to-noise will enable us to measure additional properties (e.g., star formation rate, gas phase metallicity (J. Dumayne et al., 2023)).
@@ 5 Supernova Cosmology
Following the forecasts presented in the Dark Energy Science Collaboration (DESC) Science Requirement Document (The LSST Dark Energy Science Collaboration et al., 2018), we also include an external low- sample in our cosmology analysis of 2 400 SNe Ia at . The low- SNe in our analysis follows the simulations of the Foundation SN sample (R. J. Foley et al., 2018) presented by D. O. Jones et al. (2019). The expectation, however, is that this final sample size will only be achieved towards the end of 10-years of LSST operations, even though TiDES is expected to conclude after the first five years of LSST.
@@ 5.1 Defining the SN sample
We fit all simulated SNe Ia with the SALT2 light curve fitter (J. Guy et al., 2007) based on the simulated observed photometry. To ensure meaningful light curve fits, we follow standard requirements (e.g. D. Brout et al., 2019) of: i) observations in at least two different filters with SNR5, ii) at least one data point before the time of peak brightness, , and iii) at least one data point ten days after . The first of these are already satisfied by the the spectroscopic triggering criteria in Section 3.5, and the final two are satisfied in the majority of cases due to the high cadence of WFD in LSST.
Once the model has been fit to light curves, a final subset of the SNe Ia is selected by applying the following SALT2-based cuts (e.g., M. Betoule et al., 2014; D. Brout et al., 2019): , , , days, and fit probability 0.001. Those that pass this step are considered ‘cosmologically useful’ supernovae. These selection criteria are generally adopted in SN Ia cosmology analyses to reduce contamination from highly-reddened or other peculiar thermonuclear SNe while ensuring a minimal loss of ‘cosmologically useful’ SN Ia events. We find 12 600 successful SALT2 fits in the spectroscopic sample and 131 000 successful fits in the photometric SN sample. The redshift distribution for both samples and the low- sample are shown in the top panel of Fig. 11.
@@ 5.2 SN distance estimation and bias corrections
After light-curve fitting, we measure SN Ia distances and build the redshift–distance relation (‘Hubble diagram’) from which cosmological parameters can be constrained. We estimate the SN distance modulus , which is proportional to the logarithm of the SN luminosity distance, by applying the Tripp formula (R. Tripp, 1998)
where symbols are as defined in Section 3.1.1. The nuisance parameters , and are determined following the approach presented in J. Marriner et al. (2011). We present our final Hubble diagram in Fig. 11.
The term in Equation 3 is a bias correction term that is applied to each SN to correct for survey selection effects. In this work, bias corrections are measured using the BEAMS with Bias Corrections (BBC; R. Kessler et al., 2019b) code. BBC uses large simulations of the survey to model and correct for selection effects while simultaneously fitting for the nuisance parameters and intrinsic SN scatter. For our analysis, corrections for selection effects are modelled as a function of redshift only (so-called ‘1D’ corrections) and they are estimated using the same LSST+TiDES simulated samples used to measure the cosmological constraints.
The output of BBC is a redshift-binned Hubble diagram corrected for selection effects, and the associated diagonal covariance matrix, , that includes statistical uncertainties only for each redshift bin.
@@ 5.3 Cosmological Parameters
The cosmological contours from our simulations are presented in Fig. 12 and Fig. 13. We present constraints from the LSST+TiDES SN events (both spectroscopically-confirmed SNe Ia and SNe Ia with host spec- only) and the external low- sample. We compare our results with constraints from a simulation of the Dark Energy Survey 5 year SN sample (DES-SN5YR, The DES Collaboration, 2024). The DES-SN5YR analysis constitutes the state-of-the-art of SN cosmological measurements and provides some of the tightest constraints on the Dark Energy Equation of State. The DES-SN5YR simulation used in this work is generating assuming the same cosmology used in our LSST simulations (i.e., FlatCDM cosmology with ) and reflects the statistical power and data quality of the real DES-SN5YR data. While the published DES-SN5YR include statistical and systematic uncertainties, in Fig. 12 and Fig. 13 we only consider statistical uncertainties. For a FlatCDM model, LSST+TiDES SN samples combined with external low- samples and a CMB prior (Planck Collaboration et al., 2020) provide constrains on the Dark Energy Equation of State parameter, , of 0.012, 10 times smaller compared to DES-SN5YR. For a FlatCDM model, the Figure of Merit obtained from LSST+TiDES SN samples combined with external low- samples is 85, 15 times larger than the FoM of DES-SN5YR. When fitting for a FlatCDM model, we also combine SN data with a prior based on the recent Baryonic Acoustic Oscillation (BAO) measurements published by the Dark Energy Spectroscopic Instrument (DESI) Collaboration (A. G. Adame et al., 2025) using their first year of data. The DESI-BAO-Y1 prior is applied assuming the same true cosmology used to generate the LSST and DES-SN5YR simulations. With the DESI-BAO-Y1 prior, the LSST+TiDES FoM increases to 99 (7 times larger compared to DES-SN5YR combined with the same DESI-BAO-Y1 prior).
@@ 6 Summary
We performed a cosmology analysis of the LSST+TiDES sample, including a low- sample predicted by D. O. Jones et al. (2019). We are able to constrain the dark energy equations-of-state parameter, , to a sub-2 per cent statistical-only precision, this is 10 times smaller compared to the current gold-standard (DES Collaboration et al., 2024).
@@
The DESC acknowledges ongoing support from the Institut National de Physique Nucléaire et de Physique des Particules in France; the Science & Technology Facilities Council in the United Kingdom; and the Department of Energy, the National Science Foundation, and the LSST Corporation in the United States. DESC uses resources of the IN2P3 Computing Center (CC-IN2P3–Lyon/Villeurbanne - France) funded by the Centre National de la Recherche Scientifique; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231; STFC DiRAC HPC Facilities, funded by UK BEIS National E-infrastructure capital grants; and the UK particle physics grid, supported by the GridPP Collaboration. This work was performed in part under DOE Contract DE-AC02-76SF00515. {contribution} CF led the analysis, performed the simulations, and wrote the paper. MV and MSm performed the cosmology analysis. MSu discussed the results, implications, and commented on the manuscript at all stages of the project. SFH is PI of the TiDES-RM program and wrote the relevant sections of this paper. All other co-authors are members of 4MOST/TiDES and have commented on this work.